# Relative Overfitting and Accept-Reject Framework

### v1

@@ Abstract
Currently, the scaling law of Large Language Models (LLMs) faces challenges and bottlenecks. This paper posits that noise effects, stemming from changes in the signal-to-noise ratio under diminishing marginal returns, are the root cause of these issues. To control this noise, we investigated the differences between models with performance advantages and disadvantages, introducing the concept of "relative overfitting." Based on their complementary strengths, we have proposed an application framework, Accept-Reject (AR). In Natural Language Processing (NLP), we use LLMs and Small Language Models (SLMs) as the medium for discussion. This framework enables SLMs to exert a universal positive influence on LLM decision outputs, rather than the intuitively expected negative influence. We validated our approach using self-built models based on mainstream architectures and pre-trained mainstream models across multiple datasets, including basic language modeling, long-context tasks, subject examination, and question-answering (QA) benchmarks. The results demonstrate that through our structure, compared to increasing the LLM’s parameters, we can achieve better performance improvements with significantly lower parameter and computational costs in many scenarios. These improvements are universal, stable, and effective. Furthermore, we explore the potential of "relative overfitting" and the AR framework in other machine learning domains, such as computer vision (CV) and AI for science. We hope the proposed approach can help scale laws overcome existing bottlenecks.
@@ 1 Introduction
Recently, the role of artificial intelligence in societal development has garnered increasing attention, particularly as the application of large language models (LLMs) significantly enhances human productivity. For an extended period, scaling laws 1 ; 2 ; 3 , driven by more data, greater computational power, and larger models, served as the guiding principle for training LLMs, allowing performance improvements through resource accumulation.
However, with the rapid advancement of artificial intelligence, scaling laws face increasing scrutiny 4 ; 5 , raising questions about whether they have reached a bottleneck. From the perspective of diminishing marginal returns, this observation is logical. While the debate continues over whether scaling laws have definitively hit a ceiling, evidence shows that the rate of performance improvement has notably decelerated. Furthermore, recent LLMs employing distillation techniques 6 ; 7 ; 8 ; 9 —where larger ’teacher’ models train smaller ’student’ models—demonstrate that SLMs can achieve performance comparable to LLMs (e.g., Deepseek R1 10 , Qwen 11 ). The DeepSeek R1 report specifically highlights that these distillation-based models with significantly fewer parameters perform remarkably close to their larger counterparts.
In this work, we explore an approach from the perspective of signal-to-noise ratio in model data, aiming to break the scaling law bottleneck. Data serves as the cornerstone for language models (LMs), yet along with its information, it inevitably introduces noise. Furthermore, the impact of the same data differs between a simple model exposed to limited data and a complex model utilizing vast amounts of data. Based on this premise, this paper introduces the concept of "relative overfitting" to better use the information and noise within a model. It posits that models with different structures exhibit fine-grained differences when modeling the target distribution. Within the same architecture, a model that better models the target distribution demonstrates relative overfitting compared to another model. Based on exploring this concept, we propose a universally applicable method.
Specifically, this paper proposes a universally applicable Accept-Reject (AR) framework and elaborates on it within the NLP domain, primarily using language models as vehicles. This choice is motivated by the abundance of scaling law-based model series in the field of natural language processing (NLP), which provides a robust experimental foundation. Although intuitively a model with inferior performance might seem to impede a superior one, particularly in the absence of sophisticated processing, we demonstrate that by adhering to specific proportions, a stable, positive enhancement in model performance can be achieved rather than a decline. This method, by balancing the relationship between model performance and relative overfitting, employs smaller models to influence larger ones through mechanisms of acceptance and rejection. Drawing upon the inevitable relative overfitting between models of different sizes, we utilize the Small Language Model (SLM) as a low-noise reference for the LLM. When the predictive discrepancy between them reaches a predefined threshold, we consider applying a corrective adjustment to the larger model. Compared to merely increasing the LLM’s parameter count, this structure, in many scenarios, can achieve equivalent or even superior performance enhancements for the LLM with significantly smaller parameter and computational costs. We have explored the threshold of the AR framework and posit that it could offer a potential avenue for addressing the bottlenecks in the scaling laws of large models. The method we propose is grounded in the internal relative overfitting issue within models and is not constrained by model architecture or absolute performance. It exhibits universal applicability across various model architectures and data benchmarks; that is, as long as the AR law is satisfied, even models with vastly different parameter magnitudes can be effectively influenced. Parameter magnitudes can be positively and effectively influenced.
We evaluated both smaller language models (SLMs) and larger language models (LLMs) across multiple datasets, comparing our custom-trained mainstream architectures with existing prominent language models. The results validate our method across several key aspects: its universality, shown by its applicability across all tested model architectures and datasets; in many scenarios, its effectiveness is shown by achieving better results with fewer parameters and computations, and its stability by the tendency for performance gains to increase with the main model’s size; and its flexibility, seen in offering considerably more adaptability compared to scaling up the LLM’s parameters or performing domain-specific fine-tuning. In addition, we also discussed the application of an AR framework based on the idea of relative overfitting in other areas of machine learning.
@@ 2 Related Work
The extent of actual parameter requirements has been previously investigated. Theoretically, based on the Johnson-Lindenstrauss lemma 13 , even a random projection, sampling a dimensionality reduction matrix from a Gaussian distribution, can compress N-dimensional features into $O(\log n/\epsilon^{2})$ dimensions. Experimentally, the findings of ALBERT 12 indicate that applying low-rank decomposition to BERT embeddings, reducing their dimensionality from 768 to 128, results in negligible performance degradation. This provides inspiration and evidence for our work, suggesting that when LM parameters are already substantial, there is parameter redundancy. Although increasing the parameters can enhance the granularity of the information, the associated cost of increased noise often outweighs this benefit. Furthermore, studies on word vectors, neural network interpretability, and LMs 14 ; 15 ; 16 ; 17 ; 18 ; 19 , together with research related to information and noise 20 ; 21 ; 22 ; 23 , have also informed our thinking.
Before this work, considerable research had been conducted on model fusion. Mainstream approaches include traditional ensemble methods, such as the multi-model voting schemes commonly used in Kaggle competitions 24 , which involve weighted combinations of outputs. Recent attempts within the LLM domain involve merging multiple models into a single entity 25 . Furthermore, some earlier studies have explored the fusion of output levels for LLM 26 . Unlike these studies, this article primarily explores the concept of relative overfitting and its derived application, the AR framework. Blending of models serves as a straightforward and comprehensible vehicle to articulate our ideas and validate the effectiveness of the proposed framework.
3 Relative Overfitting: What problems exist in LLM?
@@
According to the scaling law of LMs, models with more parameters generally demonstrate a greater capacity to represent details. This essentially involves a trade-off between overfitting and underfitting. For models already proven to be excellent, both overfitting and underfitting are considered well controlled. Therefore, our focus is on "relative overfitting and underfitting." A larger model possesses higher semantic granularity compared to a smaller one; that is, an LLM exhibits relative overfitting compared to an SLM. This characteristic simultaneously drives its superior performance and creates limitations in other aspects. The underlying principle is the degree of fit to the target distribution. A model that fits the target distribution well inherently exhibits relative overfitting compared to a model that fits it poorly. Even with meticulous control, relative overfitting inevitably occurs compared to a lower-performing model. Consequently, this paper explores how to influence underfitting to influence overfitting, using LLMs and SLMs merely as vehicles for discussion, given that LLMs typically demonstrate relative overfitting concerning SLMs.
@@
It is crucial to note that relative overfitting is not merely an extension of the concept of overfitting. Generally, overfitting refers to a scenario where model complexity, coupled with a small dataset, leads to excessive fitting of sample data, resulting in poor generalization to unseen datasets. However, in practice, well-tuned models following scaling laws, including LLMs, generally outperform SLMs even on unseen data. Our discussion centers on the fact that complex models possess a more detailed semantic space. Although both phenomena are due to model complexity, overfitting has associated control variables, and it can be managed by adjusting the training dataset without altering model parameters. However, relative overfitting lacks such an external control variable, making it impossible to manage this phenomenon by manipulating another quantity. This is why we emphasize the endogenous nature of relative overfitting. Simply put, overfitting depends on the relationship between model complexity and dataset size, whereas relative overfitting primarily depends on the model’s capacity to capture fine details of the target distribution.
@@
In summary, relative overfitting enables one model to identify more semantic details than another. This means that LLMs can output sentences with substantial detail. Although generally beneficial, certain drawbacks that stem from this capability inspire our method. The issue arising from this higher semantic granularity is that although the output is more detailed, there is a greater likelihood of deviating from the core subject matter. LLMs strive to match the details of the target distribution as closely as possible, which consequently increases the probability of diverging from the main topic. Conversely, SLMs, constrained by their representational capacity, cannot inherently represent extensive detail and tend to concentrate on high-frequency words. In other words, an SLM might fail to represent adjectives, focusing instead on the high co-occurrence frequency of core components, whereas an LLM will pursue the most contextually appropriate and complete expression possible. Research indicates that in most scenarios, word distributions follow Zipf’s law. Furthermore, high-frequency words are preferentially fitted within deep neural network frameworks 53 ; 54 . The improvements offered by LLMs over smaller-parameter SLMs of the same architecture are primarily concentrated on low-frequency words; the fitting difference for extremely high-frequency words is relatively small between the two. Moreover, studies suggest that within commonly used frameworks, high-frequency words tend to be overestimated, while inaccurately fitted low-frequency words are generally underestimated 55 ; 56 ; 57 ; 58 .
We define $w_{1},w_{2},\dots,w_{V}\in V$ as word vectors sorted in descending order according to the true distribution $p(w|c)$, where $V$ represents the vocabulary and $c$ is the context set for prediction.
Based on established research findings and our experiments (detailed in Appendix B.1), we propose three hypotheses:
  • Words follow Zipf’s law distribution.
  • High-frequency words are preferentially fitted during training, and the fit error increases with decreasing frequency.
  • Based on experiments (Appendix B.1), we assume that the predicted probabilities $q(w|c)$ satisfy: $q_{\text{SLM}}(w_{k}|c)\approx 0$ for $k>i$, while $q_{\text{LLM}}(w_{k}|c)\approx 0$ primarily for $k>j$, with $q_{\text{LLM}}(w_{k}|c)>0$ for some $i<k\leq j$.
We define the "core" as the high-frequency region $w_{1}$-$w_{i}$. This leads to our key theoretical result: Theorem 1. Under the proposed hypotheses, let $a\in[0,i]$, $b\in[i+1,j]$. The probability of an LLM deviating from the core region exceeds that of an SLM, formalized as:
$\frac{\sum_{a=0}^{i}|p(w_{a}|c)-q_{LLM}(w_{a}|c)|+\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{LLM}(w_{b}|c)|}{j}\geq\frac{\sum_{a=0}^{i}|p(w_{a}|c)-q_{SLM}(w_{a}|c)|}{i}$ (1)
The complete formalization of hypotheses and proofs is provided in Appendix A.
Based on the ideas discussed above, it can be articulated outside the field of NLP that there exists a relative overfitting of the larger side in scaling laws compared to the smaller side. This leads to a better modeling capability of the details of the target distribution but also increases the likelihood of deviating from the ’core’.
@@ 4 AR Framework
Based on the discussion of relative overfitting, we leverage this property to propose the ’AR framework.’ This approach utilizes the strengths of the model with a comparatively weaker fitting capability for the target distribution (in the context of relative overfitting) to influence the other model. Without loss of generality, we will continue to use SLMs and LLMs as representative models for our discussion in the NLP domain. The entire model will be introduced and discussed in four sections. In Section 4.1, we will introduce a simple application method based on the AR framework. Section 4.2 introduces the fundamental principles proposed based on the AR framework. In Section 4.3, we discuss the optimal threshold point for this framework. Finally, Section 4.4 discusses the magnitude of performance improvement achieved by the model.
@@ 4.1 Our Method
As demonstrated in the previous chapter, LLMs of the same architecture exhibit a higher probability of deviating from the core path compared to SLMs. We propose an AR framework based on this relative overfitting, specifically, a method using the SLM to influence the LLM. The underlying principle is that we balance the SLM and LLM based on the concept of relative overfitting to leverage their respective advantages, ensuring that in most situations, the SLM does not adversely affect the LLM, meaning its judgments are accepted. However, in extreme cases, the SLM can exert a rectifying effect on the LLM, leading to the rejection of the LLM’s judgment. The rationale behind this is that the existence of relative overfitting makes the LLM more prone to deviating from the core path. Therefore, if the LLM’s output significantly diverges from the SLM’s judgment, we infer that it has strayed from the correct core path. This approach allows us to preserve the LLM’s advantage in modeling detailed aspects of the target distribution under normal circumstances while exploiting the SLM’s superior advantage in modeling the core distribution (particularly concerning high-frequency words) in exceptional cases.
Based on the AR framework, we introduce a simple application training-free method. The specific implementation is as follows: utilizing our proposed AR law, we allocate proportions between the SLM and LLM, thereby identifying the AR’s threshold point. Based on this threshold point, we assign weights to the smaller and larger models and perform a weighted combination at the probability distribution layer of their outputs to obtain the final result. The weights assigned based on this threshold point (acting as a fulcrum) effectively establish a threshold to determine significant deviation. For the LLM’s output, the SLM, holding minimal weight, generally has no impact, meaning that in most situations, we primarily utilize the LLM’s detailed advantage and accept its judgment. However, once the conflict between the two reaches a critical point, that is, the SLM strongly contradicts the LLM’s judgment, we have reason to believe the LLM’s output is progressively deviating from the core path. At this juncture, we consider rejecting the LLM’s judgment. This enables us, in extreme situations, to take advantage of the core advantage of the SLM, thus achieving complementary advantages between the models (as illustrated in Figure 1).
@@
It is important to note that our emphasized viewpoint is that such discrepancies are an endogenous issue that inevitably arises from relative overfitting caused by the model structure. That is, under general circumstances, even LMs with vast performance differences possess their respective strengths and weaknesses. Our AR law allows for the effective utilization of these advantages, however slight they may seem. This forms the fundamental basis for the general and stable efficacy of our method. Naturally, in extreme cases, such as a model with exceedingly poor performance, the core region we define approximates an empty set, resulting in the SLM’s weight becoming zero, thereby exerting no influence. Specifically, applying softmax to the logit vectors $l_{i}$ ($i=1,2$) output by the two models yields:
@@
Based on the threshold point of AR, an appropriate weight $\alpha$ is determined, and a weighted sum is performed to obtain the final output:
$p_{\text{ensemble}}=\alpha\cdot p_{1}+(1-\alpha)\cdot p_{2}$ (3)
@@
Currently, when considering parallel implementation, this method incurs only linear complexity related to the vocabulary size. In many situations, compared to improving LLM performance by increasing the number of parameters, this approach offers significant savings in memory and computation time.

### v4

@@ Abstract
The scaling of Large Language Models (LLMs) currently faces significant challenges. Model assembly is widely considered a promising solution to break through these performance bottlenecks. However, current ensembling methods are primarily guided by the statistical expectation that combining multiple models over large samples will lead to performance gains. We propose an ensemble framework that transitions from such stochastic, sample-dependent methods to a regular, controllable approach based on fine-grained model segmentation. This regularity governs how models are segmented to ensure performance improvement, how the magnitude of this improvement varies with model selection, and what factors determine its theoretical maximum. To formalize this pattern, we introduce the concept of’relative overfitting,’ which is derived from the performance discrepancies between constituent models and builds a bridge between ensemble outcomes and the inherent attributes of these models. We detail the patterns of this framework within the domain of NLP and briefly describe its extensibility to other fields, such as computer vision (CV) and AI for science. Our approach was validated using both custom-built and pre-trained mainstream models across diverse benchmarks, including language modeling, long-context tasks, and question-answering (QA). The results indicate that the ensemble rules we proposed are generally effective and that we provide a rigorous proof of these rules in certain experimental scenarios. The proposed framework offers a new perspective for understanding ensemble theory and provides a systematic approach to addressing the performance bottlenecks of LLMs.
@@ 1 Introduction
The application of artificial intelligence, particularly large language models, has significantly improved people’s productivity. For an extended period, scaling laws 1 ; 2 ; 3 , driven by more data, greater computational power, and larger models, served as the guiding principle for training LLMs, allowing performance improvements through resource accumulation. However, with the rapid advancement of artificial intelligence, scaling laws face increasing scrutiny 4 ; 5 , raising questions about whether they have reached a bottleneck.
Ensemble learning has long been regarded as an effective means of overcoming the performance bottleneck of models. However, traditional ensemble learning theory relies on the improvement of performance expectations of multiple model combinations under large sample situations, which presents an irregularity; that is, we cannot ascertain whether the performance of an ensemble of two models will necessarily improve rather than decline, nor what the patterns of improvement are.
In this work, we have constructed a framework that allows performance enhancement after ensemble learning to be transformed from being irregular and dependent on large samples into a systematic ensemble achieved through a fine-grained segmentation. This enables us to better understand the operational logic of ensemble methods to enhance their effectiveness in practical applications.
At the same time, we proposed the concept of relative overfitting. It posits that models with different structures exhibit fine-grained differences when modeling the target distribution. Within the same architecture, a model that better models the target distribution demonstrates relative overfitting compared to another model. This concept can be regarded as an extension of model diversity in ensemble learning, serving as a bridge between the performance of ensemble models and the inherent properties of the individual models involved.
Specifically, this paper proposes a universally applicable Accept-Reject (AR) framework and elaborates on it within the natural language processing (NLP) domain, primarily using language models as a medium. This choice is motivated by the abundance of scaling law-based model series in the field of NLP, which provides a robust experimental foundation.
This framework introduces the concept of relative overfitting, linking the performance enhancement of ensemble models to the intrinsic properties of the models involved in the ensemble. Through a partition, we can maximize the favorable components of the ensemble while minimizing the unfavorable parts, thereby fully leveraging the advantages of each model within the ensemble. Moreover, this partitioning approach allows us to transform ensemble learning from being irregular and reliant on expectations under large samples to being systematic and controllable.
We have provided a detailed introduction to the AR law, which is based not only on our extensive experimental results across various model structures and datasets but also supported by rigorous proofs in specific scenarios. This facilitates a deeper understanding of our principles. The experimental results indicate that the principles we have proposed are universal, and by mastering the AR law, from an application perspective, this approach can achieve performance enhancements equivalent to or even superior to simply increasing the parameter count of LLMs in many cases, while significantly reducing parameter and computational costs. Additionally, from a theoretical standpoint, such regularity aids in our better understanding and control of this powerful tool, general learning.
@@ 2 Related Work
The extent of actual parameter requirements has been previously investigated. Theoretically, based on the Johnson-Lindenstrauss lemma 13 , even a random projection, sampling a dimensionality reduction matrix from a Gaussian distribution, can compress N-dimensional features into $O(\log n/\epsilon^{2})$ dimensions. Experimentally, the findings of ALBERT 12 indicate that applying low-rank decomposition to BERT embeddings, reducing their dimensionality from 768 to 128, results in negligible performance degradation. These studies inspire us to propose a relative overfitting approach as a medium to communicate the underlying patterns between the original model and the ensemble model.
Before this work, considerable research had been conducted on model fusion. Mainstream approaches include traditional ensemble methods, such as the multi-model voting schemes commonly used in Kaggle competitions 24 , which involve weighted combinations of outputs. Recent attempts within the LLM domain involve merging multiple models into a single entity 25 . Furthermore, some earlier studies have explored the fusion of output levels for LLM 26 .
Unlike these studies, our research focuses on finding a bridge between the original model and the ensemble model through the concept of relative overfitting. Within the AR framework, our aim is to transform random ensembles into regular and controllable ensembles. We demonstrate that many methods can be categorized under this framework, allowing a better application of these methods through AR law.
3 Relative Overfitting: Bridging Original and Ensemble Models
@@
According to the scaling law of LMs, models with more parameters generally demonstrate a greater capacity to represent details. This involves a trade-off between overfitting and underfitting. For models already proven excellent, both overfitting and underfitting are considered well-controlled. Therefore, our focus is on "relative overfitting and underfitting." A larger model possesses higher granularity in modeling the target distribution than a smaller one; an LLM exhibits relative overfitting compared to an SLM. This characteristic simultaneously drives its superior performance and creates limitations in other aspects. The underlying principle is the degree of fit to the target distribution. A model that fits the target distribution well inherently exhibits relative overfitting compared to a model that fits it poorly. Even with meticulous control, relative overfitting inevitably occurs compared to a lower-performing model. Consequently, this paper explores how to influence underfitting to influence overfitting, using LLMs and SLMs as the medium for discussion, given that LLMs typically relative overfitting concerning SLMs.
@@
It is crucial to note that relative overfitting is not merely an extension of the concept of overfitting. Although both phenomena are due to model complexity, overfitting has associated control variables, and it can be managed by adjusting the training dataset without altering model parameters. However, relative overfitting lacks an external control variable, making it impossible to manage this phenomenon by manipulating another quantity. This is why we emphasize the endogenous nature of relative overfitting. Simply put, overfitting depends on the relationship between model complexity and dataset size, whereas relative overfitting primarily depends on the model’s capacity to capture fine details of the target distribution.
@@
In summary, relative overfitting enables one model to identify more semantic details than another. This means that LLMs can output sentences with substantial detail. Although generally beneficial, certain drawbacks that stem from this capability inspire our method. The issue arising from this higher semantic granularity is that although the output is more detailed, there is a greater likelihood of deviating from the core subject matter. LLMs strive to match the details of the target distribution as closely as possible, consequently increasing the probability of diverging from the main topic. Conversely, SLMs, constrained by their representational capacity, cannot inherently represent extensive detail and tend to concentrate on high-frequency words. In other words, an SLM might fail to represent adjectives, focusing instead on the high co-occurrence frequency of core components, whereas an LLM will pursue the most contextually appropriate and complete expression possible. Research indicates that in most scenarios, word distributions follow Zipf’s law. Furthermore, high-frequency words preferentially fit within deep neural network frameworks 53 ; 54 . The improvements offered by LLMs over smaller-parameter SLMs of the same architecture are primarily concentrated on low-frequency words; the fitting difference for extremely high-frequency words is relatively small between the two. Moreover, studies suggest that high-frequency words tend to be overestimated within commonly used frameworks, while inaccurately fitted low-frequency words are generally underestimated 55 ; 56 ; 57 ; 58 .
We consider an LM $q$ and a Small Model (SM) $q^{\prime}$. For a specific high-frequency output class $h\in V$ (where $V$ is the set of all output classes), let $n=|S_{h}^{q}|$ and $m=|S_{h}^{q^{\prime}}|$ denote the number of instances where the LM and SM predict class $h$, respectively. Furthermore, let $P_{L}(h)$ and $P_{S}(h)$ represent the precision of LM and SM, respectively, when predicting class $h$ (i.e., $P(k_{s}=h\mid\text{model predicts }h)$).
Based on observations of model performance on high-frequency classes (detailed in Appendix B.1), we propose the following assumption.
Assumption A.2.1: For a high-frequency output class $h$, we assume:
  1. 1.
    Output Frequency Discrepancy: The SM predicts class $h$ more frequently than the LM, that is, $n<m$.
  2. 2.
    Limited LM Precision Advantage: The LM precision advantage for class $h$ is bounded, specifically satisfying:
    $|P_{L}(h)-P_{S}(h)|<\frac{(m-n)P_{S}(h)}{n}$ (1)
This assumption directly leads to our first key theoretical result concerning the overall correct predictions for such mainstay classes:
Theorem A.2.2 (Mainstay Deviation): Under the assumption, for the high-frequency output class $h$, the total number of correct predictions by the SM, $N_{S}^{h}=m\cdot P_{S}(h)$, will be greater than that of the LM, $N_{L}^{h}=n\cdot P_{L}(h)$. This is formalized as:
$N_{S}^{h}>N_{L}^{h}\quad\text{or equivalently, }\quad m\cdot P_{S}(h)>n\cdot P_{L}(h)$ (2)
The Mainstay Deviation Theorem suggests that, for common categories where an SM exhibits a higher output propensity and the LM’s precision edge is constrained, the SM may achieve a greater aggregate of correct identifications. This highlights a nuanced aspect of model scaling, where overall superiority does not always translate into superior performance in all submetric or output categories.
The complete formalization of all assumptions and detailed proofs are provided in Appendix A.
Through this concept, we can better identify patterns to segment the ensemble of the original model in detail, allowing the performance of the ensemble model to reach its optimal potential.
@@ 4 AR Framework
We proposed an AR framework that quantifies the performance of the original model in a certain scenario through the concept of relative overfitting as a basis for segmentation, enabling the ensemble model to leverage the advantages of each model while minimizing the introduction of disadvantages. Without loss of generality, we will continue to use SLMs and LLMs as representative models for our discussion in the NLP domain. Section 4.1 will introduce the AR framework based on the simplest untrained output token-level ensemble. Sections 4.2, 4.3, and 4.4 specifically introduce our ensemble theoretical principles. All code is visible: https://anonymous.4open.science/r/130.
@@ 4.1 Our Method
As mentioned above, relative overfitting allows us to use the performance differences between the original models as a basis for segmentation, where we take the better-performing model (LLM) as the primary model and the other (SLM) as the secondary model. Based on the discussion of relative overfitting, we can assume that there will always be certain positions in SLM that exert a beneficial influence on LLM, and this influence can be simply quantified by the inherent properties of the models. The concept of segmentation is to ensure that the correct corrections from SLM to LLM are as widely accepted as possible, while erroneous changes are rejected as much as possible.
We will introduce a simple ensemble method as a medium for our theoretical discussion. The specific implementation is as follows: we allocate proportions based on certain inherent properties between SLM and LLM, thereby identifying the AR’s segmentation point. Based on this segmentation point, we assign weights to the smaller and larger models and perform a weighted combination at the probability distribution layer of their outputs to obtain the final result. The weights assigned based on this segmentation point effectively establish a threshold to determine significant deviation. For the LLM’s output, the SLM, holding minimal weight, generally has no impact, meaning that in most situations, we primarily utilize the LLM’s detailed advantage and accept its judgment. However, once the conflict between the two reaches a critical point, the SLM strongly contradicts the LLM’s judgment; we have reason to believe that the LLM’s output progressively deviates from the core path. At this juncture, we consider rejecting the LLM’s judgment. This enables us, in extreme situations, to take advantage of the core advantage of the SLM, thus achieving complementary advantages between models (as illustrated in Figure 1).
This methodology was selected to elucidate the AR framework due to the inherent property that both its outputs and weights reside entirely within the probability space. This characteristic offers an intuitive basis for discussing the AR framework. The key to the AR framework lies in using relative over-collaboration as a bridge to establish the relationship between the inherent properties of the original model, which serves as the basis for finding the segmentation points. Consequently, we have also extended our discussion of the AR framework by incorporating cosine similarity metrics, as detailed in the Appendix B.3. Furthermore, based on the evidence presented in the Appendix A, we believe that more methods embodying this concept can also be incorporated within this framework to achieve a more systematic ensemble.
@@
Our segmentation framework allows us to separate adverse and beneficial influences at any fine-grained segmentation. In general, this means that even LMs with large performance disparities have their respective strengths and weaknesses. Our AR laws enable us to effectively leverage these advantages, no matter how minor they may seem. Specifically, applying softmax to the logit vectors $l_{i}$ ($i=1,2$) output by the two models yields:
@@
Based on the threshold point of AR, an appropriate weight w (Weights corresponding to the LM) is determined, and a weighted sum is performed to obtain the final output:
$p_{\text{ensemble}}=w\cdot p_{1}+(1-w)\cdot p_{2}$ (4)
@@
Currently, this method only incurs linear complexity related to vocabulary size when a parallel implementation is considered. In many situations, compared to improving LLM performance by increasing the number of parameters, this training-free method offers significant savings in memory and computation time, particularly for long sequences based on the transformer architecture.